You lend $110 today and receive a promise for repayment 10 years from now of $237. What is implied effective annual interest rate? Round your answer to 4 decimal places. For example if your answer is 3.205%, then please write down 0.0321.
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

237=110*(1+r/100)^10

(237/110)^(1/10)=(1+r/100)

(1+r/100)=1.0798

r=1.0798-1

=0.0798(Approx).